The 2nd side of a triangle is 3 inches longer than the 1st side.The 3rd side is 2 inches longer than the 2nd side. The perimeter of the triangle is 74 inches. Find the length pf the side of the triangle .
Found 2 solutions by ewatrrr, rfer:
Answer by ewatrrr(24785)


Hi,
Let x represent the length of the 1st side. The second side is (x+3) and the
third side is (x+3) + 2 Or (x + 5)
Perimeter = 74 in
x + (x+3) + (x+5) = 74 in
3x + 8 = 74 in
3x = 66in
x = 22 in, the length of the 1st side.
Second side = 25 in (22+3)
third side = 27 in (220+05)
CHECKING our Answer
22in + 25in + 27in = 74 in

Answer by rfer(16322)

x+(x+3)+((x+3)+2)=74
3x+8=74
3x=66
x=22
x+3=25
x+3+2=27